OpenStudy (anonymous):

Which is the equation of the given line? A. y = –2 B. x = 2 C. y = –2x D. x = –2

OpenStudy (anonymous):

i think B idk im 14..

OpenStudy (anonymous):

o r u in 8th

OpenStudy (anonymous):

im a freshman

OpenStudy (anonymous):

o 9th how welll im 15 and 8th i fluncked 2 grades

OpenStudy (anonymous):

its definitely b or d. lol. im a sophomore

OpenStudy (anonymous):

Life sucks man smoke weed and kick back..

OpenStudy (the_fizicx99):

^ No, just no lol.

OpenStudy (anonymous):

what the hell is wrong with you? are u high?

OpenStudy (anonymous):

ur crazy- life isn't about getting high u moron!

OpenStudy (anonymous):

it helps me focus owlet lhh!!!!!

OpenStudy (anonymous):

Its y = -2. You get this because if a line is horizontal (crosses y), has a y equation, if it is vertical (crosses x) it has an x equation. So when you know that it has a y equation, you insert the point of y (in this case, both were -2) and you have your answer: y = -2. :)

OpenStudy (anonymous):

@chucky00

OpenStudy (anonymous):

so why couldnt it be c considering the eqn of a line is y=mx + c where c could be zero??

OpenStudy (anonymous):

The equation of a line is y = mx + b, not c.

OpenStudy (anonymous):

And that equation is to find slope, not the equation of a line.

OpenStudy (anonymous):

ohh thx

OpenStudy (anonymous):

Your Welcome :)

OpenStudy (anonymous):

oh y +mx +c is the term. google said its b in the uk so for clarification 'c' is the intercept

OpenStudy (anonymous):

Um I guess. Slope-Intercept Form is always y = mx + b, no matter where you are. Google it, it may help you :) https://www.google.com/search?q=y+%3D+mx+%2B+b&rlz=1C1CHMO_enUS549US549&oq=y+%3D+mx+%2B+b&aqs=chrome..69i57j69i64.3279j0j7&sourceid=chrome&espv=210&es_sm=93&ie=UTF-8

OpenStudy (anonymous):

typo above. should be y=* not y+ but i get what your saying (Y) im new here bdw

OpenStudy (anonymous):

Nice, well welcome to OS :P If ya need any help, just mention me (@niall) :)

OpenStudy (anonymous):

ok will do thx